# Efficient Tool Use with Chain-of-Abstraction Reasoning

### v2

@@ Abstract
[…]
11footnotetext: Work done during Silin Gao’s internship at FAIR.
@@ 2 Related Work
Some other prior works focus on using LLMs for multi-step reasoning with other modules. In particular, ReAct (Yao et al., 2023b) and FireAct (Chen et al., 2023) integrate LLMs with tools into a closed loop of thought, action and observation steps. This verbose reasoning loop slows down the LLM decoding, and still incorporates tools via sequential interactions, resulting in inefficient inference. Another line of work, PAL (Gao et al., 2023) and Program of Thoughts (Chen et al., 2022) prompt LLMs to generate program-based reasoning and interact with code executors, which however heavily rely on closed source coding models, i.e., Codex (Chen et al., 2021), and are restricted to procedural arithmetic reasoning. In our work, we aim to design a more general and efficient strategy for LLMs to leverage tools, especially on multi-step reasoning scenarios.
@@ 3 Method
To construct chain-of-abstraction (CoA) data for fine-tuning LLMs, we collect question answering (QA) samples from existing open-source QA datasets (Cobbe et al., 2021; Miao et al., 2020; Yang et al., 2018), and prompt LLaMa-70B (Touvron et al., 2023a) to re-write the answer of each sampled question, as shown in Figure 2. Specifically, we prompt LLaMa-70B to label the spans in gold answers that correspond to knowledge operations (e.g., math derivations, statements based on Wikipedia references) and then to re-write the sentences with labeled spans as fillable CoA traces, where the operation results are replaced with abstract placeholders. [^4] We provide our few-shot prompting examples for CoA data re-writing in Appendix C. For example, the two derivations in the example in Figure 2 are re-written as “[$120+35=y$ 1]" and “[$290-y1=y$ 2]", respectively.
@@ 4.2 Wikipedia QA
Given a question based on Wikipedia knowledge, the model needs to first identify Wikipedia articles as references related to the question, and then reason on key knowledge in the reference articles to answer the question (as shown in the right column of Figure 1). We assume that the specialized knowledge operation in this domain is the retrieval of relevant Wikipedia articles and important named-entities, which are re-written as Wikipedia searching (WikiSearch) and named-entity recognition (NER) [^8] We use NER to extract entities from the article that bridge the former WikiSearch results to the latter WikiSearch queries. queries. Table 2 shows an example of a re-written CoA trace for Wiki QA. [^9] We include more prompting examples of Wiki QA answer re-writing in Appendix C.
@@
Model Method GSM8K ASDiv SVAMP MAWPS
AddSub SingleEQ SingleOp MultiArith All
@@
Table 4: Evaluation results on LLaMa-2 and LLaMa-2-Chat for mathematical reasoning. “All” denotes the averaged results on four MAWPS portions. Exact match rate to the final gold answer (i.e., accuracy) is reported. Best performing augmentation approach for each base model is bolded.
@@ 4.3 Baselines
We apply our CoA reasoning method to both 7B and 70B LLaMa models, and test various model versions including the first version of LLaMa (Touvron et al., 2023a) and the more advanced LLaMa-2 and LLaMa-2-Chat (Touvron et al., 2023b). We compare our method to several baselines, including: a) few-shot prompting using 8 randomly sampled QA exemplars from the original (i.e., not re-written) chain-of-thought data (CoT-FSP), b) fine-tuning with original chain-of-thought data (CoT-FT) [^12] Note that in Wiki QA domain, the HotpotQA data used for prompting or fine-tuning baselines is pre-processed to contain both gold Wikipedia articles (serving as chain-of-thought explanations) and the final answer., and c) Toolformer Schick et al. (2023) which fine-tunes LLMs on CCNet (Wenzek et al., 2020) texts augmented with API calls. For evaluation on Wiki QA, we also compared our method with FireAct (Chen et al., 2023), which fine-tunes LLMs on HotpotQA ReAct (Yao et al., 2022) trajectories distilled from GPT-4 (OpenAI, 2023).
@@ 5.1 Mathematical Reasoning
Model Method HotpotQA WQ NQ TriviaQA
Bridge Comparison Both Time
@@
Table 6: Wiki QA evaluation results on LLaMa-2-Chat-based models. “Both” denotes the overall evaluation results on both bridge and comparison portions of HotpotQA. “Time” denotes the average seconds that each agent needs to answer a question in HotpotQA. Exact match rate to the final gold answer (i.e., accuracy) is reported.
@@ 6 Conclusion
We acknowledge a few limitations in our work. First, datasets used for testing our method cannot have exhaustive coverage of all reasoning scenarios in real world. We instead consider two representative reasoning domains, i.e., mathematical reasoning and Wikipedia QA, and use English as a primary language in our testing. Furthermore, all models in our experiments use greedy decoding to generate inferences, which leaves room for applying more advanced decoding strategies, e.g., self-consistency (Wang et al., 2022) decoding. In Appendix B, we include an initial study of integrating our CoA reasoning with self-consistency decoding, where our method also achieves promising results, motivating further research on this direction. Finally, our method is tested on the setting of fine-tuning the full LLMs, which requires considerable computational resources, while more efficient model training schemes, e.g., LoRA (Hu et al., 2021), can be applied in future work.
@@ Appendix A Implementation Details
For mathematical reasoning evaluation, we extract the last number appeared in each model’s answer, and check whether the number exactly match the gold reference. The accuracy is reported as the rate of such exact match across all QAs in a test set. For Wiki QA evaluation, similar to mathematical reasoning, we extract the final answer of each model and calculate its exact match rate to the gold reference. Specifically, the final answer is supposed to be the words after “Action: finish[” for FireAct baseline, and words after “The answer is ” for other models. Our 8-shot in-domain examples used for the CoT-FSP baseline are shown in Table 13 and 14, which enables the model to provide answer with our required format for evaluation, i.e., stating its final answer after “The answer is ”. Our human evaluation on GSM8K is conducted by 5 internal domain experts from our research group. For each math question, we provide the experts with the gold answer as reference, and ask them to evaluate each model answer in anonymous manner, i.e., experts do not know which model each answer comes from. Two yes-or-no questions are asked for evaluating each model answer, including: a) whether the answer has any arithmetic error, and b) whether the answer has any reasoning error, and binary choices from the experts are collected to calculate the error rates of each model’s generation. We present our detailed instructions for human evaluation in Figure 5. Our data collection protocol is approved by our organization in terms of ethics.
@@ Appendix B Full Experimental Results
Table 7 and 8 show the full results of our experiments on math and Wiki QA domains. Our method of CoA achieves consistent improvements over baselines across various LLaMa model versions (LLaMa, LLaMa-2 and LLaMa-2-Chat), model sizes (7B and 70B), and domain benchmarks. This shows great potential of our method being generalized to new model backbones and reasoning tasks. We also present results on GSM8K subsets according to varying numbers of gold reasoning steps in Table 9, where we confirm that CoA has more robust long chain-of-thought reasoning accuracy.
@@
Self-Consistency Decoding
Besides of greedy decoding, we also test more advanced inference strategy, i.e., self-consistency (Wang et al., 2022) decoding, on our CoA reasoning method, compared to chain-of-thought baselines CoT-FSP and CoT-FT, and tool-augmented baselines Toolformer and Toolformer - Math. We test all methods on the GSM8K dataset seeded with LLaMa-2-Chat-7B. Each method samples 16 reasoning chains and uses majority voting to aggregate the 16 answers derived by the reasoning chains, to get the final answer. For the hyperparameters of sampling, we set the temperature, top-k and top-p as 1.0, 40 and 0.5, respectively. Table 10 shows our evaluation results. We find that our CoA method consistently outperforms all baseline methods when shifting from greedy decoding to self-consistency decoding. This shows that our method also has better potential to be generalized to different LLM decoding schemes.
Model Method GSM8K ASDiv SVAMP MAWPS
AddSub SingleEQ SingleOp MultiArith All
LLaMa-7B CoT-FSP 11.90 44.69 31.80 56.20 59.65 70.28 43.00 57.05
CoT-FT 30.71 53.19 42.30 55.70 69.09 77.05 54.17 64.36
CoA 35.71 56.36 51.10 67.59 80.51 85.94 68.33 75.98
LLaMa-2-7B CoT-FSP 16.38 47.85 38.40 52.41 63.39 82.03 43.33 60.53
CoT-FT 35.33 57.18 48.20 66.08 74.41 85.23 65.00 73.03
Toolformer 17.59 48.55 37.10 47.34 58.46 79.54 50.67 59.81
CoA 37.83 57.61 51.70 72.15 82.48 86.48 73.17 78.89
LLaMa-2-Chat-7B CoT-FSP 24.03 54.14 51.30 71.90 72.44 85.41 74.00 76.32
CoT-FT 35.41 59.00 46.90 58.23 72.24 85.41 73.00 73.37
CoT-FT (no ASDiv) 36.19 44.93 35.30 38.48 52.95 61.21 77.67 59.61
Toolformer 23.65 50.85 48.80 61.01 69.09 81.85 68.50 70.85
Toolformer - Math 36.01 59.18 47.60 58.99 72.44 85.94 75.50 74.43
CoA 38.29 59.57 54.20 72.41 81.89 88.26 83.00 82.13
CoA (no ASDiv) 39.73 54.19 44.40 54.18 73.62 73.49 85.33 73.27
CoA (no Tool) 35.03 58.79 51.50 68.10 74.21 86.48 77.67 77.38
LLaMa-2-Chat-70B CoT-FSP 56.18 65.94 70.60 86.08 89.17 92.88 84.50 88.23
CoT-FT 60.50 70.24 70.40 81.52 87.60 92.35 89.17 88.18
Toolformer 52.54 69.07 73.60 86.84 89.76 91.46 81.50 87.26
Toolformer - Math 61.03 70.59 73.20 85.57 91.34 91.99 92.00 90.60
CoA 62.32 71.89 73.40 86.33 94.49 93.06 92.33 91.91
GPT-J Toolformer - 40.4 29.4 - - - - 44.0
Table 7: Mathematical reasoning evaluation results.
Model Method HotpotQA WebQ. NaturalQ. TriviaQA
Bridge Comparison All
LLaMa-2-7B CoT-FSP 14.43 45.26 20.62 33.96 33.35 56.95
CoT-FT 14.85 57.36 23.39 31.50 26.93 52.32
Toolformer 14.12 42.76 20.35 37.11 34.49 57.79
CoA 22.00 57.43 29.12 34.60 38.28 58.28
LLaMa-2-Chat-7B CoT-FSP 11.69 45.46 18.47 34.65 30.91 53.48
CoT-FT 14.24 56.69 22.77 33.51 25.40 51.05
Toolformer 12.99 44.59 20.00 36.22 30.22 54.15
Toolformer - Wiki 15.68 56.42 23.86 36.61 32.96 55.08
FireAct 19.18 54.14 26.20 36.02 35.87 52.96
CoA 21.00 56.96 28.22 35.97 38.67 57.90
LLaMa-2-Chat-70B CoT-FSP 21.39 56.62 28.47 34.89 37.42 63.61
CoT-FT 23.84 63.95 31.90 34.15 39.75 62.28
Toolformer 22.24 56.09 29.04 37.16 40.42 64.31
Toolformer - Wiki 26.38 63.82 33.90 37.70 41.25 66.64
CoA 27.61 64.09 34.94 36.37 43.57 69.08
GPT-J Toolformer - - - 26.3 17.7 48.8
Table 8: Wiki QA evaluation results.
@@
Table 9: Stratified LLaMa-2-Chat-7B evaluation results on GSM8K with different gold reasoning steps. The last row reports absolute accuracy improvement of our CoA method compared to fine-tuning baseline CoT-FT.
@@
CoT-FSP 27.90
CoT-FT 39.12
Toolformer 24.56
Toolformer - Math 35.25
CoA 40.79
Table 10: Evaluation results on GSM8K with self-consistency decoding. Each model samples 16 reasoning chains with temperature 1.0, top-k 40 and top-p 0.5. All models are developed based on LLaMa-2-Chat-7B.
Appendix C Fine-Tuning Data Re-writing Details
Table 11 and 12 show the prompting examples for fine-tuning data construction of our method. We simply prompt LLaMa-70B to re-write existing math and Wiki QAs as abstract reasoning chains, which gets rid of data distillation from close-sourced LLMs, yet obtains data resources that enable more effective learning of multi-step reasoning.
Appendix D Claim of Usage
Our use of existing scientific artifacts cited in this paper is consistent with their intended use. Our developed code, data and models are intended to be used for only research purposes, any usage of our scientific artifacts that is outside of research contexts should not be allowed.

### v3

@@ Abstract
Silin Gao1,2∗, Jane Dwivedi-Yu2, Ping Yu2, Xiaoqing Ellen Tan2, Ramakanth Pasunuru2, Olga Golovneva2, Koustuv Sinha2 Asli Celikyilmaz2, Antoine Bosselut1†, Tianlu Wang2† 1EPFL, 2FAIR @ Meta 1{silin.gao,antoine.bosselut}@epfl.ch 2{silingao,janeyu,pingyu,ellenxtan}@meta.com 2{rpasunuru,olggol,koustuvs,aslic,tianluwang}@meta.com
11footnotetext: Work done during internship at FAIR.22footnotetext: Equal Supervision.
@@ 2 Related Work
Some other prior works focus on using LLMs for multi-step reasoning with other modules. In particular, ReAct (Yao et al., 2023b) and FireAct (Chen et al., 2023) integrate LLMs with tools into a closed loop of thought, action and observation steps. This verbose reasoning loop slows down the LLM decoding, and still incorporates tools via sequential interactions, resulting in inefficient inference. Another line of work, Program of Thoughts (Chen et al., 2022), DECLARATIVE (He-Yueya et al., 2023) and PAL (Gao et al., 2023) prompt LLMs to generate program-based reasoning and interact with code executors, which however heavily rely on closed source coding models, i.e., Codex (Chen et al., 2021), and are restricted to procedural arithmetic reasoning. Building on these works, CoA proposes a framework to convert natural language reasoning traces into abstract representations, and uses the abstract reasoning traces as fine-tuning data to improve tool-augmented LLMs. CoA also accelerates tool-augmented reasoning, by holistically planning the CoA traces and calling tools only once at inference time.
@@ 3 Method
To construct chain-of-abstraction (CoA) data for fine-tuning LLMs, we collect question answering (QA) samples from existing open-source QA datasets (Cobbe et al., 2021; Miao et al., 2020; Yang et al., 2018), and prompt LLaMa-70B (Touvron et al., 2023a) to re-write the answer of each sampled question, as shown in Figure 2. Specifically, we prompt LLaMa-70B to label the spans in gold answers that correspond to knowledge operations (e.g., math derivations, statements based on Wikipedia references) and then to re-write the sentences with labeled spans as fillable CoA traces, where the operation results are replaced with abstract placeholders. For example, the two derivations in the example in Figure 2 are re-written as “[$120+35=y$ 1]" and “[$290-y1=y$ 2]", respectively.
@@ 4.2 Wikipedia QA
Given a question based on Wikipedia knowledge, the model needs to first identify Wikipedia articles as references related to the question, and then reason on key knowledge in the reference articles to answer the question (as shown in the right column of Figure 1). We assume that the specialized knowledge operation in this domain is the retrieval of relevant Wikipedia articles and important named-entities, which are re-written as Wikipedia searching (WikiSearch) and named-entity recognition (NER) [^7] We use NER to extract entities from the article that bridge the former WikiSearch results to the latter WikiSearch queries. queries. Table 2 shows an example of a re-written CoA trace for Wiki QA.
@@
Model Method Use GSM8K ASDiv SVAMP MAWPS
Tool AddSub SingleEQ SingleOp MultiArith All
@@
Table 4: Evaluation results on LLaMa-2 and LLaMa-2-Chat for mathematical reasoning. “All” denotes the averaged results on four MAWPS portions. Exact match rate to the final gold answer (i.e., accuracy) is reported. For each base model, the best and second-best results are bolded and underlined, respectively. The best results labeled with are significantly better than their corresponding second-best results, with the significant test p-value $<0.05$.
@@ 4.3 Baselines
We apply our CoA reasoning method to both 7B and 70B LLaMa models, and test various model versions including the first version of LLaMa (Touvron et al., 2023a) and the more advanced LLaMa-2 and LLaMa-2-Chat (Touvron et al., 2023b). We compare our method to several baselines, including: a) few-shot prompting using 8 randomly sampled QA exemplars from the original (i.e., not re-written) chain-of-thought data (CoT-FSP), b) fine-tuning with original chain-of-thought data (CoT-FT) [^10] Note that in Wiki QA domain, the HotpotQA data used for prompting or fine-tuning baselines is pre-processed to contain both gold Wikipedia articles (serving as chain-of-thought explanations) and the final answer., and c) Toolformer Schick et al. (2023) which fine-tunes LLMs on CCNet (Wenzek et al., 2020) texts augmented with API calls. For evaluation on Wiki QA, we also compared our method with FireAct (Chen et al., 2023), which fine-tunes LLMs on HotpotQA ReAct (Yao et al., 2023b) trajectories distilled from GPT-4 (OpenAI, 2023).
@@ 5.1 Mathematical Reasoning
Method Accuracy
CoT-FSP 27.90
CoT-FT 39.12
Toolformer 24.56
Toolformer - Math 35.25
CoA 40.79
Table 6: Evaluation results on GSM8K with self-consistency decoding (seeded with LLaMa-2-Chat-7B). Each model uses majority voting to aggregate the answers of 16 sampled reasoning chains
Model Method Use HotpotQA WQ NQ TriviaQA
Tool Bridge Comparison Both Time
@@
Table 7: Wiki QA evaluation results on LLaMa-2-Chat-based models. “Both” denotes the overall evaluation results on both bridge and comparison portions of HotpotQA. “Time” denotes the average seconds that each agent needs to answer a question in HotpotQA. Exact match rate to the final gold answer (i.e., accuracy) is reported. For each base model, the best and second-best results are bolded and underlined, respectively. The best results labeled with are significantly better than their corresponding second-best results, with the significant test p-value $<0.05$.
@@
Self-Consistency Decoding
Besides of greedy decoding, we also test more advanced inference strategy, i.e., self-consistency (Wang et al., 2022) decoding, on our CoA reasoning method. We test all methods on the GSM8K dataset seeded with LLaMa-2-Chat-7B. Each method samples 16 reasoning chains and uses majority voting to aggregate the 16 answers derived by the reasoning chains, to get the final answer. For the hyperparameters of sampling, we set the temperature, top-k and top-p as 1.0, 40 and 0.5, respectively. Table 6 shows our evaluation results. We find that our CoA method consistently outperforms all baseline methods when shifting from greedy decoding to self-consistency decoding. This shows that our method also has better potential to be generalized to different LLM decoding schemes.
@@ 6 Conclusion
We acknowledge a few limitations in our work. First, datasets used for testing our method cannot have exhaustive coverage of all real-world reasoning scenarios. We instead consider two representative reasoning domains, i.e., mathematical reasoning and general open-domain (Wikipedia) QA, and use English as a primary language in our testing. Furthermore, our method is tested on the setting of fine-tuning the full LLMs, which requires considerable computational resources, while more efficient model training schemes, e.g., LoRA (Hu et al., 2021), can be applied in future work.
@@ Appendix A Implementation Details
For mathematical reasoning evaluation, we extract the last number appeared in each model’s answer, and check whether the number exactly match the gold reference. The accuracy is reported as the rate of such exact match across all QAs in a test set. For Wiki QA evaluation, similar to mathematical reasoning, we extract the final answer of each model and calculate its exact match rate to the gold reference. Specifically, the final answer is supposed to be the words after “Action: finish[” for FireAct baseline, and words after “The answer is ” for other models. Our 8-shot in-domain examples used for the CoT-FSP baseline are shown in Table 14 and 15, which enables the model to provide answer with our required format for evaluation, i.e., stating its final answer after “The answer is ”. Our human evaluation on GSM8K is conducted by 5 internal domain experts from our research group. For each math question, we provide the experts with the gold answer as reference, and ask them to evaluate each model answer in anonymous manner, i.e., experts do not know which model each answer comes from. Two yes-or-no questions are asked for evaluating each model answer, including: a) whether the answer has any arithmetic error, and b) whether the answer has any reasoning error, and binary choices from the experts are collected to calculate the error rates of each model’s generation. We present our detailed instructions for human evaluation in Figure 5. Our data collection protocol is approved by our organization in terms of ethics.
@@ Appendix B Full Experimental Results
Table 8 and 9 show the full results of our experiments on math and Wiki QA domains. Our method of CoA achieves consistent improvements over baselines across various LLaMa model versions (LLaMa, LLaMa-2 and LLaMa-2-Chat), model sizes (7B and 70B), and domain benchmarks. This shows great potential of our method being generalized to new model backbones and reasoning tasks. We also present results on GSM8K subsets according to varying numbers of gold reasoning steps in Table 10, where we confirm that CoA has more robust long chain-of-thought reasoning accuracy.
@@
More Prompting Baselines
We also compare our CoA reasoning method to more prompting-based methods PAL (Gao et al., 2023) and DECLARATIVE (He-Yueya et al., 2023), which use few-shot coding demonstrations to prompt math solutions as Python or declarative programs. Table 11 shows our comparison results on the GSM8K dataset, where all methods are seeded with LLaMa-2-Chat-7B. Without seeding with dedicated coding models (e.g., code-davinci-002), PAL and DECLARATIVE get far lower accuracy on GSM8K, which significantly under-perform our CoA method, and even ordinary CoT-FSP.
In contrast, our CoA method relies less on artificial demonstrations and distributional closeness of the seed LLM to target tasks, as CoA fine-tunes the LLM agent on pre-defined abstract reasoning chains, acquired from simple rewriting of natural language reasoning traces. Consequently, CoA is flexible in various generation formats, e.g., code and plain text, and generalizes well from mathematical reasoning to open-domain QA, which is a very different type of reasoning task. This indicates our method’s generalizability to novel reasoning schemes required by a new domain.
@@
Table 10: Stratified LLaMa-2-Chat-7B evaluation results on GSM8K with different gold reasoning steps. The last row reports absolute accuracy improvement of our CoA method compared to CoT-FT baseline.
@@
CoT-FSP 24.03
PAL 20.55
DECLARATIVE 9.86
CoA 38.29
Table 11: Comparison of CoA to prompting-based methods on GSM8K, seeded with LLaMa-2-Chat-7B.
@@
Appendix C Fine-Tuning Data Re-writing Details
Table 12 and 13 show the prompting examples for fine-tuning data construction of our method. We prompt LLaMa-70B to re-write existing math and Wiki QAs as abstract reasoning chains, which gets rid of data distillation from close-sourced LLMs, yet obtains data resources that enable more effective learning of multi-step reasoning.